Consider a particle of charge Q, moving perpendicular to a magnetic field of flux density B with a velocity v. Show that it will be deflected in a circular path, radius r, where:


r=p/BQ and p is the momentum of the charged particle.


1
Expert's answer
2021-03-14T19:21:03-0400

"F=ma\to qvB=m\frac{v^2}{r}\to qB=\frac{mv}{r}\to r=\frac{mv}{qB}=\frac{p}{qB}"


So, "r=\frac{p}{qB}" . Answer